OpenStudy (anonymous):

Can someone please help me more then what I have got... Solve by the elimination method 0.3x-0.2y=4 0.4x+0.5y=1 What is the order pair?

OpenStudy (anonymous):

helpp

OpenStudy (anonymous):

One way you can solve by elimination is by multiplying the top row by 4 and you get: 1.2x-0.8y=16 0.4x+0.5y=1 then you can recognize that in order to eliminate I variable, you must have the opposite coefficient for one of the variables to when you add the two equations it the variable ends up with a zero coeffiecient. Therefore, we can multiply the second row by a -3 and get: 1.2x-0.8y=16 -1.2x-1.5y=-3 If you add these now you will get: -2.3y=13 and y=-130/23 then you can just solve for x by substitution.

OpenStudy (anonymous):

I am really stuck on this here can you help more please. I have tried this several times and I am not getting any where

OpenStudy (anonymous):

Okay..so the first thing you must think about when asked to solve by elimination is this: "How can I manipulate one of the equations so that when it is added to another equation, it gets rid of a variable." In order to do that you must recognize common factors in the coefficients. for example, if I perform the operation .3 x 4 it equals 1.2 and if I perform the operation 0.4 x -3 is equals -1.2. 1.2 and -1.2 are now two coefficients for the variables that add to zero 1.2 + (-1.2) = 0. But when you multiply one number in an equation, you have to multiply that same number by each other coefficient in that equation, and then you can add the two, which eliminates a variable.

OpenStudy (anonymous):

Does that make sense. In other words, any equation can be multiplied by a number and as long as each coefficient in the equation is multipilied by that same number, you could say that the "meaning" of the equation does not change. By meaning I mean if you solve for a particular variable (say x), the variable will be the same. So, if I have 2x=4 you could solve for x and get x=2. Now, if you multiply that equation by a number (say 2) it looks like this: 2 [2x=4] becomes 4x=8 and solving

OpenStudy (anonymous):

Ok I think that I understand that....I learning so much from here and I am glad I just am not getting anywhere with these yet so what are the order pair.

OpenStudy (anonymous):

that gives x=2

OpenStudy (anonymous):

Ok I think that I understand that....I learning so much from here and I am glad I just am not getting anywhere with these yet so what are the order pair.

OpenStudy (anonymous):

the answer to the problem you gave is x=220/23 and y=-130/23

OpenStudy (anonymous):

thanks so I am goin to see if that is right!!

OpenStudy (anonymous):

You can check to see if it's right by plugging in the numbers and seeing if they equal 4 and 1...and they do

OpenStudy (anonymous):

So when I type them in the answer box there is still more cause it said something was wrong....Can we simplify? I tried and came up with some long decimal.

OpenStudy (anonymous):

Did you copy the problem over correctly? The answers are correct for the problem you gave me.

OpenStudy (anonymous):

Yea I did correctly...Is there infinitely many solution by any chance....I put them in (220/3,-130/23) and it said that we did something wrong cause it didnt say correct....

OpenStudy (anonymous):

There are NOT infinitely many solution. The only way there would be infinitely many solution would be if there were more unknowns than equations. For example, the following has infinitely many solutions: 4x+2y+1z=1 1x+3y+2z=4 what exactly did you enter into the answer box?

OpenStudy (anonymous):

you said that you put 220/3? That's not the correct answer for x. I said the answer is 220/23

OpenStudy (anonymous):

220/23,-130/23

OpenStudy (anonymous):

got it thanks a bunch sry

OpenStudy (anonymous):

Yep...no problem...